Ed leaves the L.A. area to visit Albuquerque, 1000 miles away, at 4:20 AM, driving at a steady speed of 70 MPH. He stops to rest every 3 hours of driving. The first rest stop is for 30 minutes and each subsequent stop is for 15 minutes more than the previous one. Find his arrival time in albuquerque to the nearest minute.
Answer by richwmiller(17219)

1000/70=100/7=14.28 hours 14 hours 17.14 minutes
so there will be 14 hours of driving
3 6 9 12
S = ½(2*30 + 15(3))4
S = 210
210 3 +30
420 6 +45
630 9 +60
840 12 +75
120 2
20 17.14
1000-840=160 miles
160-140=20 miles
20/70=2/7*60=17.14 minutes
14h 17.14
210 minutes=3 hours 30 minutes
17 hours 47.14 minutes